An airtight box has a removable lid of area 1.43E-2 m2 and negligible weight. The box is taken up a mountain where the air pressure outside the box is 8.41E+4 Pa. The inside of the box is completely evacuated. What is the magnitude of the force required to pull the lid off the box
Answer #1

2 P=8.41 x 10 N/m Area of lid= A=1.43x10-3m2 F F=PXA А = 8.41x104x1.43x10-3=120.3N ..P-